OpenStudy (anonymous):

is my answer right If the hypotenuse of an isosceles right triangle has a length of 7 cm, what is the length of one of the legs? 7√2/2

Parth (parthkohli):

\(\Large 7^{2} = 2a^{2} \) \(\Large 49 = 2a^{2} \) \(\Large 24.5 = a^2 \) \(\Large a = \sqrt{24.5} \)

OpenStudy (anonymous):

so yes it is thank u